# When LLaVA Meets Objects: Token Composition for Vision-Language-Models

### v1

@@ 1 Introduction
We evaluate the proposed framework on a suite of eight benchmark datasets, VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), ScienceQA-IMG Lu et al. (2022), POPE Li et al. (2023b), MME Fu et al. (2024), MMBench Liu et al. (2024c), and MM-Vet Yu et al. (2023). Results shows that the proposed representation is highly competitive compared to other approaches in the field. Our ablation further evaluates the scaling, as well as the token composition and the impact of pretraining on the full token set. While patch tokens alone establish a strong baseline, our experiments confirm that the additional design elements further enhance performance.
@@ 4.2 Setup
MQT Hu et al. (2024) -e2 88% 64 75.3 60.0 83.6 1464.3 63.5 67.0 51.5 28.9
Voco-LLaMa Ye et al. (2025) e2 88% 64 75.4 60.4 - - 60.5 - - -
Mask-LLaVA (ours) - e2 90% 57 74.8 60.6 83.7 1415.0 63.1 68.8 51.8 24.9
@@
MQT Hu et al. (2024) -e2 88% 36 73.7 58.8 81.9 1416.3 63.4 66.8 51.0 27.8
M3 Cai et al. (2025) -ex 75% 36 76.9 60.3 85.5 1417.2 64.8 68.2 52.8 25.4
Voco-LLaMa Ye et al. (2025) -e2 88% 32 75.3 60.2 - - 59.4 - - -
Mask-LLaVA (ours)- e2 97% 42 74.8 60.5 83.8 1402.7 63.2 68.8 52.1 25.7
MQT Hu et al. (2024)- e2 88% 16 71.1 57.6 80.8 1408.5 61.9 67.5 49.8 25.3
M3 Cai et al. (2025) -ex 75% 9 74.2 58.0 83.4 1374.8 63.1 68.0 51.9 21.7
Voco-LLaMa Ye et al. (2025) -e2 88% 16 75.4 59.4 - - 58.6 - - -
Mask-LLaVA (ours)- e2 97% 15 71.5 58.5 82.1 1395.8 62.1 68.4 52.8 21.9
@@ 4.5 Qualitative examples
We analyze qualitative results from the POPE and GQA benchmarks, comparing answers produced by Mask-LLaVA (trained with 138 tokens and tested with varying token counts) against the original LLaVA backbone (using 144 tokens at test time). Note that the examples present a random selection and were not specifically picked. For POPE, shown in Figure 3, we can observe that Mask-LLaVA performs well with respect to its robustness to hallucinations e.g. being able to spot correctly if an object is present in an image or not. Further especially on the POPE dataset, we observe that the pooling of patch tokens does not result in a loss of performance with respect to smaller fine-grained objects, which might be compensated by the respect mask embeddings. For GQA, shown in Figure 4, we can see that Mask-LLaVA further performs well with respect to the presence of the object and their attributes such as color showing, also potential to connect those attributes with location information to capture the composition of a scene. For the third example in this dataset ("Is the leather bag light brown and soft?"), Mask-LLaVA interestingly fails to provide the correct answer. This limitation may stem from CLIP embeddings (used for both the CLS token and as the basis for mask tokens) functioning additively, thus evaluating whether either condition is fulfilled rather than interpreting the question as a logical AND conjunction. More qualitative examples can be found in supplementary.

### v2

@@ 1 Introduction
We evaluate the proposed framework on a suite of eight benchmark datasets, VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), ScienceQA-IMG Lu et al. (2022), POPE Li et al. (2023b), MME Fu et al. (2024), MMBench Liu et al. (2024c), and MM-Vet Yu et al. (2023). Results show that the proposed representation is highly competitive compared to other approaches in the field. Our ablation further evaluates the scaling, as well as the token composition and the impact of pretraining on the full token set. While patch tokens alone establish a strong baseline, our experiments confirm that the additional design elements further enhance performance.
@@ 4.2 Setup
MQT Hu et al. (2024) 88% 64 75.3 60.0 83.6 1464.3 63.5 67.0 51.5 28.9
Voco-LLaMa Ye et al. (2025) 88% 64 75.4 60.4 - - 60.5 - - -
Mask-LLaVA (ours) 90% 57 74.8 60.6 83.7 1415.0 63.1 68.8 51.8 24.9
@@
MQT Hu et al. (2024) 88% 36 73.7 58.8 81.9 1416.3 63.4 66.8 51.0 27.8
M3 Cai et al. (2025) 75% 36 76.9 60.3 85.5 1417.2 64.8 68.2 52.8 25.4
Voco-LLaMa Ye et al. (2025) 88% 32 75.3 60.2 - - 59.4 - - -
Mask-LLaVA (ours) 97% 42 74.8 60.5 83.8 1402.7 63.2 68.8 52.1 25.7
MQT Hu et al. (2024) 88% 16 71.1 57.6 80.8 1408.5 61.9 67.5 49.8 25.3
M3 Cai et al. (2025) 75% 9 74.2 58.0 83.4 1374.8 63.1 68.0 51.9 21.7
Voco-LLaMa Ye et al. (2025) 88% 16 75.4 59.4 - - 58.6 - - -
Mask-LLaVA (ours) 97% 15 71.5 58.5 82.1 1395.8 62.1 68.4 52.8 21.9
@@ 4.5 Qualitative examples
We analyze qualitative results from the POPE and GQA benchmarks, comparing answers produced by Mask-LLaVA (trained with 138 tokens and tested with varying token counts) against the original LLaVA backbone (using 144 tokens at test time). Note that the examples present a random selection and were not specifically picked. For POPE, shown in Figure 3, we can observe that Mask-LLaVA performs well with respect to its robustness to hallucinations e.g. being able to spot correctly if an object is present in an image or not. Further especially on the POPE dataset, we observe that the pooling of patch tokens does not result in a loss of performance with respect to smaller fine-grained objects, which might be compensated by the respect mask embeddings. For GQA, shown in Figure 4, we can see that Mask-LLaVA further performs well with respect to the presence of the object and their attributes such as color showing, also potential to connect those attributes with location information to capture the composition of a scene. For the third example in this dataset ("Is the leather bag light brown and soft?"), Mask-LLaVA interestingly fails to provide the correct answer. This limitation may stem from CLIP embeddings (used for both the CLS token and as the basis for mask tokens) functioning additively, thus evaluating whether either condition is fulfilled rather than interpreting the question as a logical AND conjunction. More qualitative examples can be found in the appendix.
@@ 5 Conclusion
Acknowledgments
Soumya Jahagirdar is funded by the European Research Council (ERC) under the Starting Grant GraViLa (101117556). Walid Bousselham is supported by the German Federal Ministry of Education and Research (BMBF) through the project STCL (01IS22067) and by the Tübingen AI Center. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) Alvarez (2021) for funding this work by providing computing time on the GCS Supercomputer JUPITER—JUWELS at the Jülich Supercomputing Centre (JSC). This work also acknowledges support from the project ELLIOT-FM: Open Multi-Modal Foundation Models with Strong Generalization and Reasoning.